OpenStudy (ksaimouli):

3sin^2x+7sinx+2 slove for x in degress

OpenStudy (anonymous):

\[=0\]?

OpenStudy (ksaimouli):

yes

OpenStudy (anonymous):

factor

OpenStudy (ksaimouli):

(3x+2)(x+1)

OpenStudy (anonymous):

\[3z^2+7z+2=0\] \[(z+2)(3z+1)=0\]

OpenStudy (anonymous):

your factoring does not give a middle term of 7x

OpenStudy (anonymous):

\[\sin(x)+2=0\] \[\sin(x)=-2\] no chance of that happening \[3\sin(x)+1=0\] \[\sin(x)=-\frac{1}{3}\] etc

OpenStudy (ksaimouli):

ok then how 199.417 and340.529